OpenStudy (anonymous):

You put $1000 in a savings account which earns compound interest at a rate of 0.1% per month. If you don't make any deposits or withdrawls, how much will you have in your account in 3 years (36 months)? Round your answer to the nearest whole cent. A. $1036.64 B. $1331.00 C. $1003.00 D. $1430.08 @Daniellelovee

OpenStudy (danjs):

What does your book tell you about the compound interest? The probably gave you the formula right

OpenStudy (danjs):

\[Future~Value = Present~Value*(1 + rate)^{periods}\]

OpenStudy (daniellelovee):

Simple interest formula

OpenStudy (danjs):

In 36 compounding periods, the value will become... 1000(1.001)^(36)

OpenStudy (daniellelovee):

A = 1000(1 + (0.001 × 3))=?

OpenStudy (danjs):

This is compounding interest Daniell

OpenStudy (daniellelovee):

just listen to him @Tallan

OpenStudy (daniellelovee):

but how do you know that it is compound interest and not simple interest formula?

OpenStudy (danjs):

For each compounding period, the value of the account increases by 0.1%. That is the same as multiplying the value by 1.001. If you do that 36 times you get 1000*(1.001)(1.001)(1.001)......36times or Value = 1000(1.001)^(36)

OpenStudy (danjs):

You read the problem "...savings account which earns compound interest.."

OpenStudy (daniellelovee):

*facepalm*

OpenStudy (danjs):

\[Future~ Value= $1000*(1.001)^{36} \approx\]

OpenStudy (anonymous):

would it be A?